the perimeter of a rectangle is 80 inches. its length is four times as long as it's width. what is the area of the rectangle
Answer by checkley77(12844)

L=4W
2L+2W=P
2*4W+2W=80
8W+2W=80
10W=80
W=80/10
W=8 ANS. FOR THE WIDTH.
L=4*8=32 ANS FOR THE LENGTH.PROOF:
2*32+2*8=80
64+16=80
80=80
AREA=8*32=256 IN^2 IS THE AREA.